VEGAS Pulsar Project Meeting: 2015 November 9th 2:00 - 3:00pm ET

Room / Connection Details

  • GB-137 / Soc 280
  • 434-817-6443

Housekeeping Details

  • Main wiki page
  • gbsapp Mailing List
  • Location and regular date and time of meeting - agreed weekly on Mondays; GB-137 / Soc-280; 2-3pm ET
  • ETK codes:
    • Electronics: 432510.GB3015
    • Software: 893252.GB3015
    • Science: 432540.GB3015

Resource Allocations

LYNCH, Ryan 0.25


  • Vereese, Randy, Jason, Marty, Ray, Ryan, Justin, Scott, Paul

BOF file status updates

  • c1500x4096 mode
    • Ryan was able to use this with the manager on srbs-hpc1 but saw nearly 100% packet loss and extremely high load from the dedispersion thread. We will try this with newer GPU cards and on the VEGAS hardware on Thursday maintenance to see if that improves things.
    • If still seeing issues will need to sit down with Joe et al. to optimize code
      • Will need to recompile with proper architecture
    • Ryan was able to collect enough data to look at bandpass, which looks good
  • Fix for nchan calculation in coherent modes
    • Both bugs are identified and solutions are in hte c1500x4096 mode, ready for propogation to other designs
    • In the course of fixing this Randy cleaned up 18,000 flip flops from designs
    • Randy discovered some library issues but Vereese cleaned those up and pulled newest library version
      • Under Casper_11_15
      • Also incorporates ddl changes
      • Hopefully will not conflict with any existing designs, but need to keep an eye on this
  • Bandpass issues in i0800x0064, 0128, 8192
    • No updates this week
    • FFT seems to be OK, need to look for other issues
  • Duplicated signal in i0800x8192
  • Switch from 32 to to 16-bit port addresses
    • Designs now using 16-bit numbers, and this clean up is what reduced the number of flip flops
  • Use low bandwidth spectral line modes as a basis for the 100 MHz and 200 MHz pulsar mode bofs
    • Will require development on Randy's part to work out filter coefficients
    • Currently a time series is output rather than a spectrum
      • Would need to transform this and apply dedispersion filters
    • Would need to implement PFB code if it is not done on the GPU
      • Might be able to use PFB code from VEGAS
    • Could single GPU keep up with coherent dedispersion?


  • Configtool MR
    • Ryan and Justin...
  • GBT Status shared memory parameters in PSRFITS headers
    • Justin will check to see if the Manager is trying to write these parameters
    • Manager can launch the script to update the shared memory parameters when going into pulsar modes
  • Switching signal generation
    • Justin has made changes to the Manager that should invoke the relevant modules from the backend class
      • Ryan will test
    • Randy discovered that there are two values of nchan in the designs, one for switching signal generation and one for the rest of the logic
      • Switching signal needs 32 channels
      • Will need to look at switching signal phases carefully to make sure this is being handled correctly
      • Ryan can test on srbs-hpc1 and will talk to Jason and Randy about this
  • Matrix HPC
    • FITS writer
      • Writing files in lbw modes in correct format but need to test actual data
      • Justin starting on the coherent modes data path
    • Throughput tests
      • No updates (need ROACH)

GBT testing

  • Frequency shifting issue
    • Astrid SetValues fix works for testing and config tool updates should handle this naturally, so this can be considered closed.
  • Will test c1500x4096 channel mode


  • Boards and most of the parts are in
  • Waiting for quote on bandwidth limiting filters, expect to get this week
    • Order a spare, wait to see how much they cost to decide how many to order
  • Can start assembly now
  • Randy will give Vereese I/O map to begin planning, but Vereese needs to finish FLAG beamformer work before starting on interface code.
  • Most of the assembly work should be done by late next week

VEGAS/DIBAS-style backend for FAST

  • Richard, John, and Ryan talked with Karen, Mark, and Marty
    • There is interest but not much time.
    • May require delay of some VPM work, most likely with regards to the Matrix HPC. How does Ray feel about this?
  • John has put together a document detailing requirements and plan. Available via ShareLatex. (
  • Work necessar on software
    • Updates needed for calibration routines
    • Would need some work if you wanted to run spectral and pulsar modes simmultaneously
      • Could probably be done with two dealers
  • Joe sent Marty a list of work needed to bring in the latest VEGAS work
    • DIBAS FITS writer for spectral line would need updated for new keywords
    • Updates for center spike
  • Proposal to only deliver what we have working as part of normal development
    • Would need some changes to config file for different bandwidths
  • Include new HPC machine for testing as part of deal
  • They would want us to go to Beijing to use a test setup with noise sources/AP for validation
    • Telescope will not be ready so would not be traveling to telescope site
  • Software time estimate to be compiled
  • Hardware/shop time estimates to be compiled

Date / Time of Next Meeting

  • November 16

-- RyanLynch - 2015-11-09
Topic revision: r2 - 2015-11-09, RyanLynch
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback